OpenStudy (anonymous):

The potential of, A, of a charge distribution at a point on the positive x-axis is given for x in centimeters, by A= 2pi [sqrt(x^2+4)-x] A particle at x=3 is moving to the left at a rate of 0.2 cm/sec. At what rate is it potential changing?

OpenStudy (anonymous):

any help would be appreciated! Thank you!!!

OpenStudy (anonymous):

Differentiate with respect to \(t\): \[\frac{dA}{dt}=2\pi\frac{d}{dt}\left[\sqrt{x^2+4}-x\right]\] Once you do that, you plug in \(x=3\) and \(\dfrac{dx}{dt}=-0.2\) (negative because we assume the direction to the right of the origin is positive). Find \(\dfrac{dA}{dt}\).